OpenStudy (jazzyfa30):

If y varies jointly as x and z, and y = 33 when x = 9 and z = 12, find y when x = 16 and z = 22.

OpenStudy (jazzyfa30):

@PatrickJordon work it out step by step I want to see it

OpenStudy (anonymous):

Oh great thnx -_-

OpenStudy (jazzyfa30):

mmhhmm now start typing I'll start you off Y=KXZ

OpenStudy (anonymous):

ugh.

OpenStudy (jazzyfa30):

what you doing you started perfect

OpenStudy (anonymous):

then got a big number 325, then what = WRONG

OpenStudy (jazzyfa30):

okay type the equation i gave you

OpenStudy (anonymous):

Y=KXZ

OpenStudy (anonymous):

Well we do know y F***

OpenStudy (jazzyfa30):

what variables do you know substitute them in the EXACT place they are

OpenStudy (anonymous):

33=K*16*22

OpenStudy (jazzyfa30):

good know what do you do

OpenStudy (jazzyfa30):

*now

OpenStudy (anonymous):

(16(22)=352 K=352

OpenStudy (anonymous):

33= 352*16*22

OpenStudy (jazzyfa30):

do not mess with 33 so lets start over from the beginning 33=k(9)(12)

OpenStudy (anonymous):

-_-

OpenStudy (jazzyfa30):

come on you can do it.

OpenStudy (anonymous):

33=108*9*12 Doubtful

OpenStudy (jazzyfa30):

no you got 33=108 perfect but you keep forgetting the variable K.

OpenStudy (anonymous):

Ehm, isnt 108 k

OpenStudy (jazzyfa30):

no its x and z combined so 108 is xz not k

OpenStudy (anonymous):

so Divide 108 by 33

OpenStudy (anonymous):

To get, "k"

OpenStudy (jazzyfa30):

perfect yes but first let me see you equation now that you know that 108 it is xz

OpenStudy (anonymous):

33=k(108xz)

OpenStudy (jazzyfa30):

make sure your answer is a fraction once you have divided

OpenStudy (jazzyfa30):

33=108k that is what your equation should look like

OpenStudy (anonymous):

\[\frac{ 36 }{ 11 }\]

OpenStudy (jazzyfa30):

other way around

OpenStudy (anonymous):

\[\frac{ 11 }{ 36 }\]

OpenStudy (jazzyfa30):

perfect so know that we know what k is we fill it in as y= \[y=\frac{ 11 }{ 36 }xz\]what do you do next to get y

OpenStudy (anonymous):

you get rid of xz dunno

OpenStudy (jazzyfa30):

no it starts with sub

OpenStudy (anonymous):

Substitution :/

OpenStudy (anonymous):

Subtract Subway

OpenStudy (jazzyfa30):

the first one substitute the variables x and z with 9 and 12

OpenStudy (anonymous):

Ughhh.

OpenStudy (jazzyfa30):

come on you almost done

OpenStudy (anonymous):

"Almost"

OpenStudy (jazzyfa30):

this is your last step if you follow the directions i just gave youthen multiply the all up and that is your Y

OpenStudy (anonymous):

just 108,, dunno :/

OpenStudy (jazzyfa30):

okay \[y=\frac{ 11 }{ 36 }*16*12\] you should get a fraction

OpenStudy (jazzyfa30):

and what does Subway have to do with math

OpenStudy (anonymous):

i was trying to be funny :P

OpenStudy (jazzyfa30):

lol

OpenStudy (anonymous):

\[\frac{ 176 }{ 3 }\] How the H

OpenStudy (jazzyfa30):

sorry had to do something for mom and you should have gotten 968/9

OpenStudy (anonymous):

blah.... ok

OpenStudy (jazzyfa30):

don't get smart with me

OpenStudy (anonymous):

I'm pretty sure i cant b smart with any one xp

OpenStudy (anonymous):

brb

OpenStudy (jazzyfa30):

anyways

OpenStudy (anonymous):

back